# Flattening Hierarchies with Policy Bootstrapping

### v1

@@ 1 Introduction
Goal-conditioned reinforcement learning (GCRL) specifies tasks by desired outcomes, alleviating the burden of defining reward functions over the state-space and enabling the training of general policies capable of achieving a wide range of goals. Offline GCRL extends this paradigm to leverage existing datasets of reward-free trajectories and has been likened to the simple self-supervised objectives that have been successful in training foundation models for other areas of machine learning [49, 31]. However, the conceptual simplicity of GCRL belies practical challenges in learning accurate value functions and, consequently, effective policies for goals requiring complex, long-horizon behaviors. These limitations call into question its applicability as a general and scalable objective for learning foundation policies [5, 34, 18] that can be efficiently adapted to a diverse array of control tasks.
@@
Hierarchical reinforcement learning (HRL) is commonly used to address these challenges and is particularly well-suited to the recursive subgoal structure of goal-reaching tasks, where reaching distant goals entails first passing through intermediate subgoal states. Goal-conditioned HRL exploits this structure by learning a hierarchy composed of multiple levels: one or more high-level policies, tasked with generating intermediate subgoals between the current state and the goal; and a low-level actor, which operates over the primitive action space to achieve the assigned subgoals. These approaches have achieved state-of-the-art results in both online [28, 24] and offline GCRL [33], and are especially effective in long-horizon tasks. However, despite the strong empirical performance of HRL, it suffers from major limitations as a scalable pretraining strategy. In particular, the modularity of hierarchical policy architectures, fixed to specific levels of temporal abstraction, precludes unified task representations and necessitates learning a generative model over the subgoal space to interface between policy levels.
@@
Learning to predict intermediate goals in a space that may be as high-dimensional as the raw observations poses a difficult generative modeling problem. To ensure that subgoals are physically realistic and reachable in the allotted time, previous work often implements additional processing and verification of proposed subgoals [53, 15, 7, 50]. An alternative is to instead predict in a compact learned latent subgoal space, but simultaneously optimizing subgoal representations and policies results in a nonstationary input distribution to the low-level actor, which can slow and destabilize training [43, 24]. The choice of objective for learning such representations, ranging from autoregressive prediction [38, 51] to metric learning [42, 30, 26], remains an open question and adds significant complexity to the design and tuning of hierarchical methods.
@@
Following the tantalizing promise that flat, one-step policies can be optimal in fully observable, Markovian settings [36], this work aims to isolate the core advantages of hierarchies for offline GCRL and distill them into a simpler training recipe for a single, unified policy. We begin our empirical analysis by revisiting a state-of-the-art hierarchical method for offline GCRL that significantly outperforms previous approaches on a range of long-horizon goal-reaching tasks. Beyond the original explanation based on improved value function signal-to-noise ratio, we find that separately training a low-level policy on nearby subgoals improves sampling efficiency. We reframe this hierarchical approach as a form of implicit test-time bootstrapping on subgoal-conditioned policies, revealing a conceptual connection to earlier methods that learn subgoal generators and bootstrap directly from subgoal-conditioned policies to train a flat, unified goal-conditioned policy.
@@
Building on these insights, we present an inference-based theoretical framework that unifies these ideas and yields Subgoal Advantage-Weighted Policy Bootstrapping (SAW), a novel policy extraction objective for offline GCRL. SAW uses advantage-weighted importance sampling to bootstrap on subgoals sampled directly from data, capturing the long-horizon strengths of hierarchies in a single, flat policy without requiring a generative subgoal model. In evaluations across 20 state- and pixel-based offline GCRL datasets, our method matches or surpasses all baselines in diverse locomotion and manipulation tasks and scales especially well to complex, long-horizon tasks, being the only existing approach to achieve nontrivial success in the humanoidmaze-giant environment.
@@ 2 Related Work
Our work builds on a rich body of literature encompassing goal-conditioned RL [19], offline RL [21, 23], and hierarchical RL [8, 41, 39, 4, 43]. The generality of the GCRL formulation enables powerful self-supervised training strategies such as hindsight relabeling [2, 12] and state occupancy matching [25]. These are often combined with approaches that exploit the recursive subgoal structure of GCRL: either implicitly via quasimetric learning [47], probabilistic interpretations [16, 54], and contrastive learning [10, 55]; or explicitly through hierarchical decomposition into subtasks [28, 24, 13, 33]. Despite these advances, learning remains difficult for distant goals due to sparse rewards and discounting over time. Many methods use the key insight that actions which are effective for reaching an intermediate subgoal between the current state and the goal are also effective for reaching the final goal. Such subgoals are typically selected via planning [17, 52, 14], searching within the replay buffer [9], or, most commonly, sampling from generative models. Hierarchical methods in particular generate subgoals during inference and use them to query “subpolicies” trained on shorter horizon goals, which are generally easier to learn [40, 3]. Our method also leverages the ease of training subpolicies to effectively learn long-horizon behaviors, but aims to learn a flat, unified policy while avoiding the complexity of training generative models to synthesize new subgoals.
Policy bootstrapping. Our work is most closely related to Reinforcement learning with Imagined Subgoals [6, RIS], which, to our knowledge, is the only prior work that performs bootstrapping on policies, albeit in the online setting. Similar to goal-conditioned hierarchies, RIS learns a generative model to synthesize “imagined” subgoals that lie between the current state and the goal. Unlike HRL approaches, however, it regresses the full-goal-conditioned policy towards the subgoal-conditioned target, treating the latter as a prior to guide learning and exploration [Figure 1]. While RIS yields a flat policy for inference, it still requires the full complexity of a hierarchical policy, including a generative model over the goal space. In contrast, our work extends the core benefits of subgoal-based bootstrapping to offline GCRL with an advantage-based importance weight on subgoals sampled from dataset trajectories, eliminating the need for a subgoal generator altogether.
@@ 3 Preliminaries
Offline value learning: We use a goal-conditioned, action-free variant of implicit Q-learning [20, IQL] referred to as goal-conditioned implicit value learning [33, GCIVL]. The original IQL formulation modifies standard value iteration for offline RL by replacing the $\max$ operator with an expectile regression, in order to avoid value overestimation for out-of-distribution actions. GCIVL replaces the state-action value function with a value-only estimator
$\min_{\psi}\mathcal{L}_{\mathrm{GCIVL}}(\psi)=\mathbb{E}_{s,a,g\sim p^{\mathcal{D}}}\left[\ell_{\tau}^{2}\left(r(s,g)+\gamma\bar{V}\left(s^{\prime},g\right)-V(s,g)\right)\right],$ (2)
where $\ell_{\tau}^{2}(x)=|\tau-\mathbbm{1}(x<0)|x^{2}$ is the expectile loss parameterized by $\tau\in[0.5,1)$ and $\bar{V}(\cdot)$ denotes a target value function. Note that GCIVL is optimistically biased in stochastic environments, since it directly regresses towards high-value transitions without using Q-values to marginalize over action-independent stochasticity.
Offline policy extraction: To learn a target subpolicy, we use Advantage-Weighted Regression [35, AWR] to extracts a policy from a learned value function. AWR reweights state-action pairs according to their exponentiated advantage with an inverse temperature hyperparameter $\alpha$, via the objective
$\max_{\pi}\mathcal{J}_{\mathrm{AWR}}(\pi)=\mathbb{E}_{s,a,g\sim\mathcal{D}}\left[e^{\alpha(Q(s,a,g)-V(s,g))}\log\pi(a\mid s,g)\right],$ (3)
thus remaining within the support of the data without requiring an additional behavior cloning penalty.
@@ 4 Understanding Hierarchies in Offline GCRL
In this section, we seek to identify the core reasons behind the empirical success of hierarchies in offline GCRL that can be used to guide the design of a simpler training objective for a flat policy. We first review previous explanations for the benefits of HRL and propose an initial algorithm that seeks to capture these benefits in a flat policy, but find that it still fails to close the performance gap to Hierarchical Implicit Q-Learning [33, HIQL], a state-of-the art method. We then identify an additional practical benefit of hierarchical training schemes and show how HIQL exploits this from a policy bootstrapping perspective.
@@ 4.1 Hierarchies in online and offline GCRL
Previous investigations into the benefits of hierarchical RL in the online setting attribute their success to improved exploration [39] and training value functions with multi-step rewards [29]. They demonstrate that augmenting non-hierarchical agents in this manner can largely close the performance gap to hierarchical policies. However, the superior performance of hierarchical methods in the offline GCRL setting, where there is no exploration, calls this conventional wisdom into question.
HIQL is a state-of-the-art hierarchical offline GCRL method that extracts a high-level policy over subgoals and a low-level policy over primitive actions from a single goal-conditioned value function trained using standard one-step temporal difference learning. HIQL achieves significant performance gains across a number of complex, long-horizon navigation tasks purely through improvement on the policy extraction side, without needing multi-step rewards to train the value function as done in Nachum et al. [29]. While this does not preclude the potential benefits of multi-step rewards for offline GCRL, it does demonstrate that the advantages of hierarchies are not limited to temporally extended value learning, in line with previous claims that the primary bottleneck in offline RL is policy extraction and not value learning [32].
@@ 4.2 Value signal-to-noise ratio in offline GCRL
Instead, HIQL addresses a separate “signal-to-noise ratio” (SNR) issue in value functions conditioned on distant goals, where a combination of sparse rewards and discounting makes it nearly impossible to accurately determine the advantage of one primitive action over another with respect to distant goals. By separating policy extraction into two levels, the low-level actor can instead evaluate the relative advantage of actions with respect to nearby subgoals and the high-level policy can utilize multi-step advantage estimates to get a clearer learning signal with respect to distant goals.
@@
$\tilde{A}(s_{t},a_{t},g)\approx A\left(s_{t},a_{t},w\right),$ (4)
where $w\sim\pi^{h}_{\text{sg}}(w\mid s_{t},g)$ is a subgoal sampled from a high-level policy $\pi^{h}$, and the sg subscript indicates a stop-gradient operator. Apart from using this advantage to directly train a flat policy with AWR, we use the same architectures, sampling distributions, and training objective for $\pi^{h}$ as HIQL. Despite large gains over one-step policy learning objectives in several navigation tasks, GCWAE still underperforms its hierarchical counterpart, achieving a 55% success rate on antmaze-large-navigate compared to 90% for HIQL without subgoal representations and 16% for GCIVL with AWR.
@@ 4.3 It’s easier to find good (dataset) actions for closer goals
While diagnosing this discrepancy, we observed that training statistics for the two methods were largely identical except for a striking difference in the mean action advantage $A(s_{t},a_{t},w)$. The advantage was significantly lower for GCWAE, which samples “imagined” subgoals $w\sim\pi^{h}(\cdot\mid s,g)$ from a high-level policy, than HIQL, which samples directly from the $k$-step future state distribution $w\sim p^{\mathcal{D}}(s_{t+k}\mid s_{t})$ of the dataset. This leads us to an obvious but important insight: in most cases, dataset actions are simply better with respect to subgoals sampled from nearby future states in the trajectory than to distant goals or “imagined” subgoals generated by a high-level policy. The dataset is far more likely to contain high-advantage actions for goals sampled at the ends of short subsequences, whereas optimal state-action pairs for more distant goals along the trajectory are much rarer due to the combinatorial explosion of possible goal states as the goal-sampling horizon increases.
The practical benefits of being able to easily sample high-advantage state-action-goal tuples are hinted at in Park et al. [31], who pose the question “Why can’t we use random goals when training policies?” after finding that offline GCRL algorithms empirically perform better when only sampling (policy) goals from future states in the same trajectory as the initial state. While their comparison focuses on in-trajectory versus random goals instead of nearer versus farther in-trajectory goals, we hypothesize both observations are driven by similar explanations.

### v3

@@ 1 Introduction
Goal-conditioned reinforcement learning (GCRL) specifies tasks by desired outcomes, alleviating the burden of defining reward functions over the state-space and enabling the training of general policies capable of achieving a wide range of goals. Offline GCRL extends this paradigm to existing datasets of reward-free trajectories, and has been likened to the simple self-supervised objectives that have been successful in training foundation models for other areas of machine learning [yang_what_2023, park_ogbench_2024]. However, the conceptual simplicity of offline GCRL belies practical challenges in learning accurate value functions and, consequently, effective policies for goals requiring complex, long-horizon behaviors. These limitations call into question its applicability as a general and scalable objective for learning foundation policies [black_pi_0_2024, park_foundation_2024, intelligence_pi_05_2025] that can be efficiently adapted to diverse control tasks.
@@
Hierarchical reinforcement learning (HRL) is commonly used to address these challenges and is particularly well-suited to the recursive subgoal structure of goal-reaching tasks, where reaching distant goals entails first passing through intermediate subgoal states. Goal-conditioned HRL exploits this structure by learning a hierarchy composed of multiple levels: one or more high-level policies, tasked with generating intermediate subgoals between the current state and the goal; and a low-level actor, which operates over the primitive action space to achieve the assigned subgoals. These approaches have achieved state-of-the-art results in both online [levy_learning_2018, nachum_data-efficient_2018] and offline GCRL [park_hiql_2023], and are especially effective in long-horizon tasks with sparse rewards.
@@
However, despite the strong empirical performance of HRL, it suffers from major limitations as a scalable pretraining strategy. In particular, the modularity of hierarchical policy architectures, fixed to specific levels of temporal abstraction, precludes unified task representations and necessitates learning a generative model over the subgoal space to interface between policy levels. Learning to predict intermediate goals in a space that may be as high-dimensional as the raw observations poses a difficult generative modeling problem. To ensure that subgoals are physically realistic and reachable in the allotted time, previous work often implements additional processing and verification of proposed subgoals [zhang_generating_2020, czechowski_subgoal_2021, hatch_ghil-glue_2024, zawalski_fast_2022]. An alternative is to instead predict in a compact learned latent subgoal space, but simultaneously optimizing subgoal representations and policies results in a nonstationary input distribution to the low-level actor, which can slow and destabilize training [vezhnevets_feudal_2017, levy_learning_2018]. The choice of objective for learning such representations, ranging from autoregressive prediction [seo_reinforcement_2022, zeng_goal-conditioned_2023] to temporal metric learning [tian_model-based_2020, nair_r3m_2023, ma_vip_2023], remains an open question and adds significant complexity to the design and tuning of hierarchical methods.
@@
Following the tantalizing promise that flat, one-step policies can be optimal in fully observable, Markovian settings [puterman_markov], this work aims to isolate the core advantages of hierarchies for offline GCRL and distill them into a training recipe for a single, unified one-step policy. We first conduct an empirical analysis on a state-of-the-art hierarchical method for offline GCRL that significantly outperforms previous approaches on a range of long-horizon goal-reaching tasks. Beyond the original explanation based on improved value function signal-to-noise ratio, we find that separately training a low-level policy on nearby subgoals improves sampling efficiency. We then reframe hierarchical policies as a form of implicit test-time bootstrapping on subgoal-conditioned policies, revealing a theoretical connection to earlier methods that learn subgoal generators and bootstrap directly from subgoal-conditioned policies to train a flat, unified goal-conditioned policy.
@@
Building on these insights, we construct an inference problem over optimal subgoal states that unifies these ideas and yields Subgoal Advantage-Weighted Policy Bootstrapping (SAW), a novel policy extraction objective for offline GCRL. SAW uses advantage-weighted importance sampling to bootstrap on in-trajectory waypoint states, capturing the long-horizon strengths of hierarchies in a single, flat policy without requiring a generative subgoal model. In evaluations across 20 state- and pixel-based offline GCRL datasets, our method matches or surpasses all baselines in diverse locomotion and manipulation tasks and scales especially well to complex, long-horizon tasks, being the only existing approach to achieve nontrivial success in the humanoidmaze-giant environment.
@@ 2 Related Work
Our work builds on a rich body of literature encompassing goal-conditioned RL [kaelbling_learning_1993], offline RL [lange_batch_2012, levine_offline_2020], and hierarchical RL [dayan_feudal_1992, sutton_between_1999, jong_utility_2008, bacon_option-critic_2017, vezhnevets_feudal_2017]. The generality of the GCRL formulation enables powerful self-supervised training strategies such as hindsight relabeling [andrychowicz_hindsight_2017, ghosh_learning_2020] and state occupancy matching [ma_offline_2022, eysenbach_contrastive_2022, zheng_contrastive_2023, sikchi_score_2023]. These are often combined with approaches that exploit the recursive subgoal structure of GCRL: either implicitly by enforcing quasimetric structure on the goal-conditioned value function [wang_optimal_2023], or explicitly through hierarchical decomposition into subgoals [nachum_data-efficient_2018, levy_learning_2018, gupta_relay_2020, park_hiql_2023]. Despite these advances, learning remains difficult for distant goals due to sparse rewards and discounting over time. Many methods rely on the key insight that actions which are effective for reaching an intermediate subgoal between the current state and the goal are also effective for reaching the final goal. Such subgoals are typically selected via planning [huang_mapping_2019, zhang_world_2021, hafner_deep_2022], searching within the replay buffer [eysenbach_search_2019], or, most commonly, sampling from generative models. Hierarchical methods in particular generate subgoals during inference and use them to query “subpolicies” trained on shorter-horizon goals, which are generally easier to learn [strehl_reinforcement_2009, azar_minimax_2017]. Our method also leverages the ease of training subpolicies to effectively learn long-horizon behaviors, but aims to learn a flat, unified policy while avoiding the complexity of training generative models to synthesize new subgoals.
Policy bootstrapping: Our work is most closely related to Reinforcement learning with Imagined Subgoals [chane-sane_goal-conditioned_2021, RIS], which, to our knowledge, is the only prior work that performs bootstrapping on policies, albeit in the online setting. Similar to goal-conditioned hierarchies, RIS learns a generative model to synthesize “imagined” subgoals that lie between the current state and the goal. Unlike HRL approaches, however, it regresses the full-goal-conditioned policy towards the subgoal-conditioned target, treating the latter as a prior to guide learning and exploration [Figure 1]. While RIS yields a flat policy for inference, it still requires the full complexity of a hierarchical policy, including a generative model over the goal space. In contrast, our work extends the core benefits of subgoal-based bootstrapping to offline GCRL with an advantage-based importance weight on subgoals sampled from dataset trajectories, eliminating the need for a subgoal generator altogether.
@@ 3 Preliminaries
Offline value learning: We use a goal-conditioned, action-free variant of implicit Q-learning [kostrikov_offline_2021, IQL] referred to as goal-conditioned implicit value learning [park_hiql_2023, GCIVL]. The original IQL formulation modifies standard value iteration for offline RL by replacing the $\max$ operator with an expectile regression, in order to avoid value overestimation for out-of-distribution actions. GCIVL replaces the critic $Q(s,a,g)$ with an action-free estimator $\bar{Q}(s,a,g)=r(s,g)+\gamma\bar{V}\left(s^{\prime},g\right)$, minimizing
$\mathcal{L}_{\mathrm{GCIVL}}(\psi)=\mathbb{E}_{s,a,g\sim p^{\mathcal{D}}}\left[\ell_{\tau}^{2}\left(r(s,g)+\gamma\bar{V}\left(s^{\prime},g\right)-V_{\psi}(s,g)\right)\right],$ (2)
where $\ell_{\tau}^{2}(x)=|\tau-\mathbbm{1}(x<0)|x^{2}$ is the expectile loss parameterized by $\tau\in[0.5,1)$ and $\bar{V}(\cdot)$ denotes a target value function. Note that using the action-free estimator with expectile regression is optimistically biased in stochastic environments, since it directly regresses towards high-value transitions without using Q-values to marginalize over environment stochasticity.
Offline policy extraction: To learn a target subpolicy, we use Advantage-Weighted Regression [peng_advantage-weighted_2019, AWR] to extract a policy from a learned value function. AWR reweights state-action pairs according to their exponentiated advantage with an inverse temperature hyperparameter $\alpha$, maximizing
$\mathcal{J}_{\mathrm{AWR}}(\pi)=\mathbb{E}_{s,a,g\sim\mathcal{D}}\left[e^{\alpha(\bar{Q}(s,a,g)-V(s,g))}\log\pi(a\mid s,g)\right]$ (3)
and thus remaining within the support of the data without requiring an additional behavior cloning penalty. As with GCIVL, we use the action-free estimate $\bar{Q}(s,a,g)$ to compute the advantage.
@@ 4 Understanding Hierarchies in Offline GCRL
In this section, we seek to identify the core reasons behind the empirical success of hierarchies in offline GCRL that can be used to guide the design of a training objective for a simpler, flat policy. We first review previous explanations for the benefits of HRL and propose an initial algorithm that seeks to capture these benefits in a flat policy, but find that it still fails to close the performance gap to a state-of-the art method. We then identify an additional practical benefit of hierarchical training schemes and show how HIQL exploits this from a policy bootstrapping perspective.
@@ 4.1 Hierarchies in online and offline GCRL
Previous investigations into the benefits of hierarchical RL in the online setting attribute their success to improved exploration [jong_utility_2008] and training value functions with multi-step rewards [nachum_why_2019]. They demonstrate that augmenting non-hierarchical agents in this manner can largely close the performance gap to hierarchical policies. However, the superior performance of hierarchical methods in the offline GCRL setting, where there is no exploration, calls this conventional wisdom into question.
Our empirical investigation focuses on Hierarchical Implicit Q-Learning [park_hiql_2023, HIQL], a simple yet effective hierarchical method for offline GCRL that learns a high-level policy over subgoals and a low-level policy over primitive actions from a single goal-conditioned value function [Figure 1]
$J_{\pi^{h}}\left(\theta_{h}\right)$ $=\mathbb{E}_{(s_{t},s_{t+k})\sim\mathcal{D},g\sim p(g)}\left[\exp\left(\beta^{h}\cdot\tilde{A}^{h}\left(s_{t},s_{t+k},g\right)\right)\log\pi_{\theta_{h}}^{h}\left(s_{t+k}\mid s_{t},g\right)\right]$
$J_{\pi^{\ell}}\left(\theta_{\ell}\right)$ $=\mathbb{E}_{\left(s_{t},a_{t},s_{t+1},s_{t+k}\right)\sim\mathcal{D}}\left[\exp\left(\beta^{\ell}\cdot\tilde{A}^{\ell}\left(s_{t},a_{t},s_{t+k}\right)\right)\log\pi_{\theta_{\ell}}^{\ell}\left(a_{t}\mid s_{t},s_{t+k}\right)\right],$ (4)
where $\beta^{h}$ and $\beta^{\ell}$ are inverse temperature hyperparameters and $\tilde{A}$ is an approximation of the advantage function described further in Appendix F.1. Importantly, the value function is trained with standard one-step temporal-difference (TD) learning (GCIVL) instead of multi-step rewards [nachum_why_2019], isolating HIQL’s significant performance gains across a number of complex, long-horizon navigation tasks to improvements in policy extraction. While this does not preclude the potential benefits of multi-step rewards for offline GCRL, it does demonstrate that the advantages of hierarchies are not limited to temporally extended value learning, in line with previous claims that the primary bottleneck in offline RL is policy extraction and not value learning [park_is_2024].
@@ 4.2 Value signal-to-noise ratio in offline GCRL
Instead, HIQL claims to address a separate “signal-to-noise ratio” (SNR) issue in goal-conditioned value functions when the goal is very far away, where a combination of sparse rewards and discounting makes it nearly impossible to accurately determine the advantage of one primitive action over another with respect to distant goals. This is also known as the action gap phenomenon [farahmand_action-gap_2011]. By separating policy extraction into two levels, the low-level actor can instead evaluate the relative advantage of actions for reaching nearby subgoals and the high-level policy can utilize multi-step advantage estimates to get a better notion of progress towards distant goals.
@@
$A^{w}(s_{t},a_{t},g)\approx A\left(s_{t},a_{t},\pi^{h}_{\text{sg}}(w\mid s_{t},g)\right),$ (5)
where the sg subscript indicates a stop-gradient operator. Apart from using this advantage to extract a flat policy with AWR, we use the same architectures, sampling distributions, and training objective for $\pi^{h}$ as HIQL. Despite large gains over one-step policy learning objectives in several navigation tasks, GCWAE still underperforms its hierarchical counterpart, achieving a 75% success rate on antmaze-large-navigate compared to 90% for HIQL without subgoal representations and 16% for GCIVL with AWR policy extraction [Appendix H].
@@ 4.3 It’s easier to find good (dataset) actions for closer goals
While diagnosing this discrepancy, we observed that training statistics for the two methods were largely identical except for a striking difference in the mean action advantage $A(s_{t},a_{t},w)$. The advantage was significantly lower for GCWAE, which samples “imagined” subgoals $w\sim\pi^{h}(\cdot\mid s,g)$ from a high-level policy to train the flat policy; than HIQL, which samples from the real $k$-step future state distribution $w\sim p^{\mathcal{D}}(s_{t+k}\mid s_{t})$ of the dataset to train the low-level policy (and then imagines subgoals during inference). This leads us to an obvious but important insight: in most cases, dataset actions are simply better with respect to subgoals sampled from nearby future states in the trajectory than to distant goals or “imagined” subgoals generated by a high-level policy. The dataset is far more likely to contain high-advantage actions for goals sampled at the ends of short subsequences, whereas optimal state-action pairs for more distant goals along the trajectory are much rarer due to the combinatorial explosion of possible goal states as the goal-sampling horizon increases.
The practical benefits of being able to easily sample high-advantage state-action-goal tuples are hinted at in park_ogbench_2024, who pose the question “Why can’t we use random goals when training policies?” after finding that offline GCRL algorithms empirically perform better when only sampling (policy) goals from future states in the same trajectory as the initial state. While their comparison focuses on in-trajectory versus random goals instead of nearer versus farther in-trajectory goals, we hypothesize both observations are driven by similar explanations: namely, the solutions reached by constrained policy optimization algorithms are dependent on the base sampling distribution over actions [levine_offline_2020, park_is_2024].